OpenStudy (anonymous):

Solve the system of equations. x - 4y = 5 x = 3y - 2 (__,__)

jhonyy9 (jhonyy9):

x-4y=5 x-3y=-2 multiplie by -1 and get -x+3y=2 so than assume the same terms x-4y=5 -x+3y=2 ---------- 0 -y=7 y = -7 x-4(-7)=5 x+28=5 x=5-28 x=-23 x=-23 y=-7 hope this will help you